AAVSOnet news 2013-10-29
Here is the current status of the AAVSOnet telescopes.
BSM_Berry After cleaning the CCD, some defects are seen. We are investigating this; may have to swap out cameras.
BSM_Hamren closed down for a few weeks until the site operator returns.
BSM_NM closed down until ~Nov9 as operator is out of town.
BSM_South working fine
W30 limping along; new mount will be installed in a few weeks
SRO working fine. backlog of images to process for researchers.
OC61 working fine, making improvements to pipeline
TMO61 working fine. backlog of images to process, getting pipeline implemented.